Zandile is about to purchase a car for R191 810,00 on terms from a car dealer. The terms
are R23 235 deposit and the rest paid in equal quarterly payments of R13 292,48 at the end
of each quarter for four years. The interest rate is 11,5% per year, compounded quarterly.
Considering the amortisation schedule, the principal repaid during the third quarter of the
first year, is equal to
1
Expert's answer
2020-04-14T13:04:08-0400

The question does not describe whether interest on deposit or a percentage of depreciation are given.

"FV=Pv(1+\frac{i}{m})^{mn}=23235*(1+\frac{11.5}{4})^{16}=23235(1.02875)^{16}=36 567.92"


"\frac{36567.92}{16}=2285.495"